how do you determine the equation and graph the line with a slope of 2/3, containing the point (6,-5)

Answer by mananth(16946)

m= 2/ 3

Plug value of the slope and point ( 6 , -5 ) in
Y = m x + b
-5.00 = 4 + b
b= -5 - 4
b= -9
So the equation will be
Y = 2/3x -9
y= 2/ 3 x + -9

Assume values of x and plug in the equation to get values of y.
Plot the points
x 0 6 3 -3
y= -9 -5 -7 -11
(x,y) (0,-9),(6,-5),(3,-7),(-3,-11)



m.ananth@hotmail.ca